# Lyapunov exponents and geodesic stability of Schwarzschild black hole in the non-commutative gauge theory of gravity

### v3

@@ IV Conclusions
In this study, we examined the geodesic motion of both massive and massless test particles in the NC Schwarzschild spacetime. Our approach was based on modifying the Schwarzschild BH metric using the NC gauge theory of gravity abdellah1 . We computed corrections to the effective potential up to second order in $\Theta$ and obtained modifications to both the affine parameter and the coordinate time in the NC Schwarzschild BH. Our results indicate that due to the influence of non-commutativity, both massive and massless particles require an infinite amount of time to reach either the NC singularity or the NC event horizon.
Furthermore, we analyzed the stability of geodesic motion by utilizing the Lyapunov exponent to assess the stability of circular orbits for both massive and massless particles. For massive particles, we determined the angular frequency $\hat{\Omega}_{c}$, the time period $\hat{T}_{c}$, and both the proper-time and coordinate-time Lyapunov exponents, $\hat{\lambda}_{p}$ and $\hat{\lambda}_{c}$, respectively. Notably, the proper-time Lyapunov exponent revealed the presence of a stable orbital region between two unstable regions near the event horizon, which aligns with previous findings abdellah1 . For massless particles, our analysis of the effective potential uncovered a novel type of motion near the event horizon, indicating the presence of a stable circular orbit. As a result, a second photon sphere emerges in NC spacetime. Consequently, the NC Schwarzschild BH features two distinct photon spheres: a stable inner photon sphere and an unstable outer photon sphere. By employing the coordinate-time Lyapunov exponent $\hat{\lambda}_{c}^{\text{Null}}$, we confirmed the instability of the outer photon sphere, while the inner photon sphere remains stable—a feature unique to NC spacetime. Additionally, we computed the critical exponent $\hat{\gamma}^{\text{Null}}$ to further evaluate the instability of null circular orbits in the NC Schwarzschild BH.
Lastly, we explored the impact of non-commutativity on the BH shadow. Our results suggest a resemblance between the effects of mass and non-commutativity, implying that non-commutativity strengthens the gravitational field of the BH. Key insights from our analysis include the connection between newly identified stable circular orbits near the event horizon and non-commutative effects, as well as an estimated upper bound on the NC parameter, which is approximately $\Theta^{\text{Phy}}\sim 10^{-32}$ m.

### v4

@@ IV Conclusions
In this study, we examined the geodesic motion of both massive and massless test particles in the NC Schwarzschild spacetime. Our approach was based on modifying the Schwarzschild BH metric using the NC gauge theory of gravity abdellah1 . In this geometry, we show that $\hat{g}_{tt}\hat{g}_{rr}\neq 1$ is a necessary condition to satisfy the four energy conditions. This implies that the BH exhibits a double-photon-sphere structure: the inner photon sphere, which is stable, and the outer one, which is unstable. We computed corrections to the effective potential up to second order in $\Theta$ and obtained modifications to both the affine parameter and the coordinate time in the NC Schwarzschild BH. Our results indicate that due to non-commutativity, both massive and massless particles require an infinite amount of time to reach either the NC singularity or the NC event horizon.
Furthermore, we analyzed the stability of geodesic motion using the Lyapunov exponent to assess circular orbit stability for both massive and massless particles. For massive particles, we determined the angular frequency $\hat{\Omega}_{c}$, the time period $\hat{T}_{c}$, and both the proper-time and coordinate-time Lyapunov exponents, $\hat{\lambda}_{p}$ and $\hat{\lambda}_{c}$, respectively. Notably, the proper-time Lyapunov exponent revealed a stable orbital region between two unstable regions near the event horizon, aligning with previous findings abdellah1 . For massless particles, our analysis of the effective potential uncovered a novel type of motion near the event horizon, indicating the presence of a stable circular orbit. As a result, a second photon sphere emerges in NC spacetime. Consequently, the NC Schwarzschild BH features two distinct photon spheres: a stable inner photon sphere and an unstable outer photon sphere. Using the coordinate-time Lyapunov exponent $\hat{\lambda}_{c}^{\text{Null}}$, we confirmed the instability of the outer photon sphere, while the inner photon sphere remains stable— a feature unique to NC spacetime. Additionally, we computed the critical exponent $\hat{\gamma}^{\text{Null}}$ to further assess the instability of null circular orbits in the NC Schwarzschild BH.
Lastly, we explored the impact of non-commutativity on the BH shadow. Our results suggest that non-commutativity strengthens the BH’s gravitational field, mimicking the effect of increased mass. Key insights from our analysis include the connection between newly identified stable circular orbits near the event horizon and non-commutative effects, as well as an estimated upper bound on the NC parameter through the shadow radius, unstable photon sphere, and outer ISCO for massive test particles, which is approximately $\Theta^{\text{Phy}}\sim(10^{-33}-10^{-34})$ m. Furthermore, our results highlight the impact of this geometry on orbital stability and geodesic motion around deformed compact objects (such as BHs), predicting new types of motion near the event horizon and introducing a new fundamental length scale, supporting the emergence of quantum effects at macroscopic levels. The study of other astronomical phenomena could further enhance our understanding of quantum gravity effects on large-scale systems.